# Real enumerative invariants relative to the toric boundary and their refinement

### v1

@@ 4.3 Wall-crossing
(5) Suppose that $\xi$ is as in Lemma 4.2(2iii) and ${\Phi}(p)=w_{i}^{\sigma}(t^{*})=:w$. Observe that $({\Phi}\big{|}_{\widehat{C}_{s}})^{*}(w)=2l_{s}p$, $s=1,2$, where $l_{1}+l_{2}=k_{i}^{\sigma}$. Suppose that ${\Phi}\big{|}_{\widehat{C}_{1}},{\Phi}\big{|}_{\widehat{C}_{2}}$ both are immersions and that $l_{1}\leq l_{2}$.
Define the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$, $0<\zeta\ll 1$, by picking one mobile point in $C_{1}\cap\widehat{\boldsymbol{w}}(t^{*})\setminus\{w_{i}^{\sigma}(t^{*})\}$ and the other in $C_{2}\cap\widehat{\boldsymbol{w}}(t^{*})\setminus\{w_{i}^{\sigma}(t^{*})\}$. It defines a one-dimensional subvariety of ¯$\overline{\mathcal{M}}_{0,n}(\Delta)$, and let =${\mathcal{F}}=\{[{\Phi}_{t}:(\widehat{C}^{(t)},{\boldsymbol{p}}_{t})\to{\operatorname{Tor}}(P_{\Delta})]\}_{|t-t^{*}|<\zeta}$ be one of the irreducible germs of that variety at $\xi$, where $\widehat{C}^{(t)}\simeq{\mathbb{P}}^{1}$ as $t\neq 0$. The one-dimensional family of curves $C^{(t)}={\Phi}_{t}(\widehat{C}^{(t)})\in|{\mathcal{L}}_{P_{\Delta}}|$, $|t-t^{*}|<\zeta$, has a tangent cone at $C_{1}\cup C_{2}$ spanned by $C_{1}\cup C_{2}$ and some curve $C^{*}\in|{\mathcal{L}}_{P_{\Delta}}|\setminus\{C_{1}\cup C_{2}\}$.
@@
$\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C_{1})}\delta(C_{1},z)+c_{1}({\mathcal{L}}_{P_{\Delta}})[C_{1}]-1=C_{1}^{2}+1\ .$
@@
where h.o.t. designates the terms above the Newton diagram, and
@@
Consider the tropical limit at $t_{1}\to 0$. It includes a subdivision of the triangle $T={\operatorname{conv}}\{(0,1),(4l,0),(02)\}$ and certain limit curves in the toric surfaces associated with the pieces of the subdivision. Since locally $C^{(t)}$ is an immersed cylinder, we obtain that the union of the limit curves is a curve of arithmetic genus zero, which finally allows the only following tropical limit: the triangle $T$ is the unique piece of the subdivision, and
@@
is rational. It also follows, that ${\mathcal{F}}$ is smooth, regularly parameterized by $t_{1}$ (i.e., $\lambda=1$), and its real part submersively projects onto the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$. By the patchworking theorem [21, Theorems 3.1 and 4.2], we uniquely restore ${\mathcal{F}}$ as long as we compute the coefficients $a_{r1}^{0}$, $20\leq r\leq 2l_{1}-$ 2 (here $a_{2l-1,1}^{0}$ is determined by $C^{*}$). The rationality of the curve (20) can be expressed as follows. Write equation (20) in the form $y^{2}-2P(x)y+\eta_{1}\eta_{2}x^{4l}=0$, then resolve with respect to $y$:
@@
where $\lambda_{1},\lambda_{2},\lambda_{3}$ can be computed out of $\eta_{1},\eta_{2}$, and $a_{2l-1,1}^{0}$. As noticed in [22, Proof of Proposition 6.1], there are exactly two real solutions, one corresponding to a curve with the real part, shown in Figure 4(c) by lines inside the four Newton triangles designating four real quadrants, and this curve has one hyperbolic node and $22l-$ 2 non-real nodes, while the other curve, shown in Figure 4(d), has $12l-$ 1 elliptic nodes. Bringing the complex orientations to the play, we see that the former solution appears when the local real branches $({\mathbb{R}}C_{1},w)$ and ${\mathbb{R}}C_{2},w)$ are cooriented, while the latter solution corresponds to the case when these real branches have opposite orientations, see Figures 4(c,d). That is, on both sides of the path $\{{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$ we encounter four elements of $\overrightarrow{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))$ having quantum invariants
@@
It remains to notice that the Welschinger signs (8) of the elements with quantum indices $\kappa_{1}+\kappa_{2},-\kappa_{1}-\kappa_{2}$ are the same on both sides, since the modified limit curves shown in Figure 4(c) for $t>t^{*}$ and $t<t^{*}$ are obtained from each other by reflection with respect to the vertical axis. The same holds for the elements with quantum indices $\kappa_{1}-\kappa_{2},\kappa_{2}-\kappa_{1}$ (see Figure 4(d)). This completes the proof of the constancy of the numbers $W_{0}^{\kappa}(t)$ in the considered case $l_{1}=l_{2}$.
@@
(see Figure 5(a)), while the monomials on the segment $[(0,1),(2l_{1},1)]$ sum up to $\eta_{1}(x+\lambda t_{1})^{2l_{1}}y$, where $\lambda\neq 0$ is uniquely determined by the coefficient $a_{2l_{1}-1,1}^{0}$ coming from $C^{*}$. The two limit curves have branches intersecting each other with multiplicity $2l_{1}$ along the toric divisor ${\operatorname{Tor}}([(0,1),(2l_{1},0)])$ (see Figure 5(b)). The genuine geometry of $C^{(t)}$ can be recovered when we deform that intersection point into $12l_{1}-$ 1 nodes. The modification describes such a deformation as a replacement of the intersection point by one of the $2l_{1}$ modified limit curves, among which exactly two are real, and their real parts are shown in Figure 5(c,d) (cf. Figures 4(c,d)): one of them has a hyperbolic node and $22l_{1}-$ 2 non-real nodes, and the other $12l_{1}-$ 1 elliptic nodes. As in the preceding paragraph, the former modified limit curve fits the case when the local real branches $({\mathbb{R}}C_{1},w)$, $({\mathbb{R}}C_{2},w)$ are cooriented, which means that on each side of the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$ we have two elements of $\overrightarrow{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))$ with quantum indices $\kappa_{1}+\kappa_{2},-\kappa_{1}-\kappa_{2}$, while all four elements have the same Welschinger sign (8) since the constructions for $t>t^{*}$ and $t<t^{*}$ are symmetric with respect to the vertical axis. Hence, the constancy of the numbers $W_{0}^{\kappa}(t)$. The same holds in the case of the opposite orientation of the branches $({\mathbb{R}}C_{1},w)$, $({\mathbb{R}}C_{2},w)$ with the use of the second modified limit curve. We also point out that the final deformation including the modification still depends on $t_{1}=t-t^{*}$.
@@
The remaining case is as follows: ${\Phi}:\widehat{C}_{1}\to{\operatorname{Tor}}(P_{\Delta})$ is an $2l_{1}$-multiple covering of a line through $w$ and ${\Phi}:\widehat{C}_{2}\to{\operatorname{Tor}}(\Sigma)$ is an immersion onto a rational curve having a smooth branch at $w$ intersecting ${\operatorname{Tor}}(\sigma)$ with multiplicity $2l_{2}$. Here we remove the restriction $l_{1}\leq l_{2}$. We extend the range of the coordinate system $x,y$ to a neighborhood of the line $C_{1}={\Phi}(\widehat{C}_{1})$, assuming that the axis $\{x=0\}$ is the whole line $C_{1}$. Then the Newton diagram of ${\Phi}_{*}(\widehat{C})$ in this neighborhood of $C_{1}$ is as shown in Figure 6(a) by fat lines, where $s=C_{1}C_{2}$. Following the recipe used in the preceding cases, we obtain that the tropical limit of the family $\{C^{(t)}\}_{|t-t^{*}|<\zeta}$ as $t_{1}:=t-t^{*}\to 0$ includes the triangle ${\operatorname{conv}}\{(0,1),(2l_{1},1),(2(l_{1}+l_{2}),0)\}$ and the rectangle ${\operatorname{conv}}\{(0,1),(0,s+1),(2l_{1},1),(2l_{1},s+1)\}$ (see Figure 6(b)), and, in addition, the part of the equation of $C^{(t)}$ restricted to the union of these polygons is as follows:
$y\cdot((x-\lambda_{1}t_{1})^{2l_{1}}-a(x-\lambda_{2}t_{1})^{2l_{1}}y)\prod_{j=1}^{s-1}(y-y_{j})-\eta x^{2(l_{1}+l_{2})}=0\ ,$
where $y_{1},...,y_{s-1}$ are the coordinates of the points $C_{1}\cap C_{2}\setminus\{w\}$, $\lambda_{1}\neq 0$ is determined by $C^{*}$, $\lambda_{2}t_{1}\neq 0$ is the current coordinate of the mobile point from $C_{1}\cap\widehat{\boldsymbol{w}}(t^{*})\setminus\{w\}$, and $a,\eta>0$ are determined by $C_{2}$. Again the two limit curves have intersection of multiplicity $2l_{1}$ in a point on the toric divisor ${\operatorname{Tor}}([(0,1),(2l_{1},1)])$. The genuine geometry of $C^{(t)}$ in a neighborhood of $w$ is obtained by deforming this tangency point by means of the two real modified limit curves, which we used in the preceding paragraph. Then, the argument from the preceding paragraph in the same manner completes the proof of the constancy of the numbers $W_{0}^{\kappa}(t)$ in the considered wall-crossing bifurcation.
(5b) The fact that the (global) family ${\mathcal{F}}$ is smooth, realizes the local wall crossing described in part (5a), and transversally crosses the wall amounts by the standard argument used in part (4), to the condition (19). The latter condition can be verified exactly as in part (4).
(6) Suppose that $\xi$ is as in Lemma 4.2(2iv). Without loss of generality we can suppose that
@@
and that ${\Phi}$ is ramified at the points of $\widehat{\boldsymbol{w}}(t^{*})$ on the toric divisors associated with the segments $[(0,2m),(2p,0)]$ and $[(0,2m),(2q,0)]$, while the two points $w_{1}^{\sigma}(t),w_{2}^{\sigma}(t)\in{\boldsymbol{w}}(t)$ on the toric divisor ${\operatorname{Tor}}(\sigma)$, where $\sigma=[(2p,0),(2q,0)]$, merge to one point as $t\to t^{*}$. Then $C={\Phi}({\mathbb{P}}^{1})$ admits a parametrization
@@
This deformation is regularly parameterized by the difference of the $x$-coordinates of $w_{2}^{\sigma}(t)$ and $w_{1}^{\sigma}(t)$ equal to $2ma\lambda+O(\lambda^{2})$. It follows from formulas (21) that the corresponding element $\xi(t)\in\overline{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))$ has the same orientations at a point on the toric divisor ${\operatorname{Tor}}([(0,2m),(2p,0)])$ for $\lambda<0$ and for $\lambda>0$, and so does for the point on the toric divisor ${\operatorname{Tor}}([(0,2m),(2q,0)])$, while the local branches at $w_{1}^{\sigma}(t)$, $w_{2}^{\sigma}(t)$ have opposite orientations with respect to the orientation of ${\operatorname{Tor}}_{\mathbb{R}}(\sigma)$. Furthermore, $\xi(t)$ has exactly two elliptic nodes in a neighborhood of each elliptic node of $C$. Thus, the constancy of $W_{0}^{\kappa}(t)$ follows. $\Box$
[…]
Figure 6: Proof of Lemma 4.3, part (5), III
@@
Take two sequences $\widehat{\boldsymbol{w}}(0)=(\widehat{\boldsymbol{w}}_{\partial}(0),w_{0}(0))$ and $\widehat{\boldsymbol{w}}(1)=(\widehat{\boldsymbol{w}}_{\partial}(1),w_{0}(1))$, satisfying the conditions of Theorem 3.7. We may assume that $\widehat{\boldsymbol{w}}_{\partial}(0),\widehat{\boldsymbol{w}}_{\partial}(1)\in\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ for some $\tau>0$. Then, we join these sequences by a generic path $\{\widehat{\boldsymbol{w}}_{\partial}(t)\}_{0\leq t\leq 1}$, in $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ and the points $w_{0}(0)$, $w_{0}(1)$ by a generic path $\{w_{0}(t)\}_{0\leq t\leq 1}$ in ${\mathcal{Q}}$, and verify the constancy of $W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ in all possible wall-crossing events, for all $\kappa$.

### v2

@@ 4.3 Wall-crossing
(5) Suppose that $\xi$ is as in Lemma 4.2(2iii) and ${\Phi}(p)=w_{i}^{\sigma}(t^{*})=:w$. Observe that $({\Phi}\big{|}_{\widehat{C}_{r}})^{*}(w)=2l_{r}p$, $r=1,2$, where $l_{1}+l_{2}=k_{i}^{\sigma}$.
(5a) Suppose that ${\Phi}\big{|}_{\widehat{C}_{1}},{\Phi}\big{|}_{\widehat{C}_{2}}$ both are immersions and that $l_{1}\leq l_{2}$.
Define the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$, $0<\zeta\ll 1$, by picking one mobile point in $C_{1}\cap\widehat{\boldsymbol{w}}(t^{*})\setminus\{w\}$ and the other in $C_{2}\cap\widehat{\boldsymbol{w}}(t^{*})\setminus\{w\}$. It defines a one-dimensional subvariety of ¯$\overline{\mathcal{M}}_{0,n}(\Delta)$, and let ${\mathcal{V}}=\{[{\Phi}_{t}:(\widehat{C}^{(t)},{\boldsymbol{p}}_{t})\to{\operatorname{Tor}}(P_{\Delta})]\}_{|t-t^{*}|<\zeta}$ be one of the irreducible germs of that variety at $\xi$, where $\widehat{C}^{(t)}\simeq{\mathbb{P}}^{1}$ as $t\neq 0$. The one-dimensional family of curves $C^{(t)}={\Phi}_{t}(\widehat{C}^{(t)})\in|{\mathcal{L}}_{P_{\Delta}}|$, $|t-t^{*}|<\zeta$, has a tangent cone at $C_{1}\cup C_{2}$ spanned by $C_{1}\cup C_{2}$ and some curve $C^{*}\in|{\mathcal{L}}_{P_{\Delta}}|\setminus\{C_{1}\cup C_{2}\}$.
@@
$\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C_{1})}\delta(C_{1},z)-C_{1}K_{{\operatorname{Tor}}(P_{\Delta})}-1=C_{1}^{2}+1\ .$
@@
where h.o.t. denotes the terms above the Newton diagram, and
@@
Consider the tropical limit at $t_{1}\to 0$. It includes a subdivision of the triangle $T={\operatorname{conv}}\{(0,1),(4l,0),(0,2)\}$ and certain limit curves in the toric surfaces associated with the pieces of the subdivision. Since locally $C^{(t)}$ is an immersed cylinder, we obtain that the union of the limit curves is a curve of arithmetic genus zero, which finally allows the only following tropical limit: the triangle $T$ is the unique piece of the subdivision, and
@@
is rational. It also follows, that ${\mathcal{V}}$ is smooth, regularly parameterized by $t_{1}$ (i.e., $\lambda=1$), and its real part submersively projects onto the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$. By the patchworking theorem [27, Theorems 3.1 and 4.2], we uniquely restore ${\mathcal{V}}$ as long as we compute the coefficients $a_{r1}^{0}$, $20\leq r\leq 2l_{1}-$ 2 (here $a_{2l-1,1}^{0}$ is determined by $C^{*}$). The rationality of the curve (23) can be expressed as follows. Write equation (23) in the form $y^{2}-2P(x)y+\eta_{1}\eta_{2}x^{4l}=0$, then resolve with respect to $y$:
@@
where $\lambda_{1},\lambda_{2},\lambda_{3}$ can be computed out of $\eta_{1},\eta_{2}$, and $a_{2l-1,1}^{0}$. As noticed in [28, Proof of Proposition 6.1], there are exactly two real solutions, one corresponding to a curve with the real part, shown in Figure 3(c) by lines inside the four Newton triangles designating four real quadrants, and this curve has one hyperbolic node and $22l-$ 2 non-real nodes, while the other curve, shown in Figure 3(d), has $12l-$ 1 elliptic nodes. Bringing the complex orientations to the play, we see that the former solution appears when the local real branches $({\mathbb{R}}C_{1},w)$ and ${\mathbb{R}}C_{2},w)$ are oriented coherently, while the latter solution corresponds to the case when these real branches have opposite orientations, see Figures 3(c,d). That is, on both sides of the path $\{{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$ we encounter four elements of $\overrightarrow{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))$ having respective quantum indices
@@
It remains to notice that the Welschinger signs (9) and (10) of the elements with quantum index $\kappa_{1}+\kappa_{2}$ or $-\kappa_{1}-\kappa_{2}$ are the same on both sides, since the modified limit curves shown in Figure 3(c) for $t>t^{*}$ and $t<t^{*}$ are obtained from each other by reflection with respect to the vertical axis. The same holds for the elements with quantum index $\kappa_{1}-\kappa_{2}$ or $\kappa_{2}-\kappa_{1}$ (see Figure 3(d)). This completes the proof of the constancy of the functions $W_{0}^{\kappa}(t)$ and $\widetilde{W}_{0}^{\kappa}(t)$ in the considered case $l_{1}=l_{2}$.
@@
(see Figure 4(a)), while the monomials on the segment $[(0,1),(2l_{1},1)]$ sum up to $\eta_{1}(x+\lambda t_{1})^{2l_{1}}y$, where $\lambda\neq 0$ is uniquely determined by the coefficient $a_{2l_{1}-1,1}^{0}$ coming from $C^{*}$. The two limit curves have branches intersecting each other with multiplicity $2l_{1}$ along the toric divisor ${\operatorname{Tor}}([(0,1),(2l_{1},0)])$, see Figure 4(b). The genuine geometry of $C^{(t)}$ can be recovered when we deform that intersection point into $12l_{1}-$ 1 nodes. The modification describes such a deformation as a replacement of the intersection point by one of the $2l_{1}$ modified limit curves, among which exactly two are real, and their real parts are shown in Figure 4(c,d) (cf. Figures 3(c,d)): one of them has a hyperbolic node and $22l_{1}-$ 2 non-real nodes, and the other has $12l_{1}-$ 1 elliptic nodes. As in the preceding paragraph, the former modified limit curve fits the case when the local real branches $({\mathbb{R}}C_{1},w)$, $({\mathbb{R}}C_{2},w)$ are oriented coherently, which means that on each side of the path $\{\widehat{\boldsymbol{w}}(t)\}_{|t-t^{*}|<\zeta}$ we have two elements of $\overrightarrow{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))$ with quantum indices $\kappa_{1}+\kappa_{2},-\kappa_{1}-\kappa_{2}$, while all four elements have the same Welschinger sign (9) or (10) since the constructions for $t>t^{*}$ and $t<t^{*}$ are symmetric with respect to the vertical axis. Hence, we obtain the constancy of the functions $W_{0}^{\kappa}(t)$, $\widetilde{W}_{0}^{\kappa}(t)$. The same holds in the case of the opposite orientation of the branches $({\mathbb{R}}C_{1},w)$, $({\mathbb{R}}C_{2},w)$ with the use of the second modified limit curve. We also point out that the final deformation including the modification still depends on $t_{1}=t-t^{*}$.
@@
The realizability of the deformations described in this item by variation of the constraints requires an independent variation of the coefficients of the monomials under the Newton diagram $\Gamma$ while keeping the total $\delta$-invariant outside the point $w$ and keeping the given intersection number $2k_{i^{\prime}}^{\sigma^{\prime}}$ with the toric divisor ${\operatorname{Tor}}(\sigma^{\prime})$ at a point close to $w_{i^{\prime}}^{\sigma^{\prime}}$, for each $(i^{\prime},\sigma^{\prime})\neq(i,\sigma)$. This can be expressed as
$h^{1}(C,{\mathcal{J}}_{Z/C}\otimes_{{\mathcal{O}}_{C}}{\operatorname{in}}^{*}{\mathcal{L}}_{P_{\Delta}})=0,$ (24)
where $C=C_{1}\cup C_{2}$, $C_{r}={\Phi}(\widehat{C}_{r})$, $r=1,2$, the zero-dimensional scheme $Z\subset C$ is supported at ${\operatorname{Sing}}\hskip 1.70709pt(C)\cup\widehat{\boldsymbol{w}}(t^{*})$ and defined by the ideal $J^{cond}(C,z)$ at each $z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}$, by the ideal $\{\varphi\in{\mathcal{O}}_{C,w_{i^{\prime}}^{\sigma^{\prime}}}\ :\ (\varphi\cdot C)_{w_{i^{\prime}}^{\sigma^{\prime}}}\geq 2k_{i^{\prime}}^{\sigma^{\prime}}-1\}$ at each $w_{i^{\prime}}^{\sigma^{\prime}}\in\widehat{\boldsymbol{w}}(t^{*})\setminus\{w\}$, and by the ideal generated by the monomials on the Newton diagram $\Gamma$ and above it at the point $w$. This $h^{1}$-vanishing lifts to the disjoin union of $\widehat{C}_{1}$ and $\widehat{C}_{2}$ in the form
$h^{1}(\widehat{C}_{r},{\mathcal{O}}_{\widehat{C}_{r}}({\boldsymbol{d}}_{r}))=0,\quad r=1,2,$ (25)
where
$\deg{\boldsymbol{d}}_{r}=C^{2}_{r}-(C^{2}_{r}+C_{r}K_{{\operatorname{Tor}}(P_{\Delta})}+2)-(-C_{r}K_{{\operatorname{Tor}}(P_{\Delta})}-|C_{r}\cap{\operatorname{Tor}}(\partial P_{\Delta})|+1)$
$=|C_{r}\cap{\operatorname{Tor}}(\partial P_{\Delta})|-3\geq-1>-2,\quad r=1,2,$ (26)
which certifies (25).
@@
(5b) The remaining case is as follows: ${\Phi}:\widehat{C}_{1}\to{\operatorname{Tor}}(P_{\Delta})$ is an $2l_{1}$-multiple covering of a line $C_{1}\simeq{\mathbb{P}}^{1}$ through $w$ and ${\Phi}:\widehat{C}_{2}\to{\operatorname{Tor}}(P_{\Delta})$ is a (birational) immersion onto a rational curve $C_{2}$ having a smooth branch at $w$ intersecting ${\operatorname{Tor}}(\sigma)$ with multiplicity $2l_{2}$. Here, we remove the restriction $l_{1}\leq l_{2}$. We extend the range of the coordinate system $x,y$ to a neighborhood of the line $C_{1}$, assuming that $C_{1}=\{x=0\}$. Then, the Newton diagram of ${\Phi}_{*}(\widehat{C})$ in this neighborhood of $C_{1}$ is as shown in Figure 5(a) by fat lines, where $N$ is the intersection multiplicity of the curves $C_{1},C_{2}$. Following the recipe used in the preceding cases and taking into account that the curves $C^{(t)}$, $0<|t-t^{*}|\ll 1$, are rational, we obtain that
  • the combinatorial part of the tropical limit consists of the triangle ${\operatorname{conv}}\{(0,1),(2l_{1},1),(2(l_{1}+l_{2}),0)\}$ and the rectangle ${\operatorname{conv}}\{(0,1),(0,N),(2l_{1},1),(2l_{1},N)\}$ (see Figure 5(b));
  • the limit curve associated with the triangle is a rational curvewhich intersect the toric divisor ${\operatorname{Tor}}([(0,1),(2l_{1},1)])$ at one point, where it is smooth, while the limit curve associated with the rectangle is split into $2l_{1}$-multiple line intersecting the two horizontal toric divisors and $N-1$ distinct lines intersecting the two vertical toric divisors; here, the position of $N-1$ horizontal lines is determined by the intersection points of $C_{2}$ with ${\operatorname{Tor}}([(2l_{1},1),(2l_{1},N)])$, and the position of the vertical line is determined by the intersection point of he limit curve from the preceding item with ${\operatorname{Tor}}([(0,1),(2l_{1},1)])$.
Note that the $2l_{1}$-multiple component of the limit curve associated with the rectangle is, in fact, the image of a $2l_{1}$-multiple covering ${\Phi}:\widehat{C}_{1}\to C_{1}$, and the point $p$ is mapped to the point $x=\tau$ on the toric divisor ${\operatorname{Tor}}([(0,1),(2l_{1},1)])$. The deformation can be written in the form
[…]
where $\tau=t-t^{*}$, the non-zero constants $a,y_{1},...,y_{N-1}$ are determined by $C_{2}$, and $f(x,y)$ is the sum of monomials strictly right to the Newton diagram with coefficients determined by $C_{2}$. The higher order terms (h.o.t.) yield the appearance of $2l_{1}(N-1)$ nodes of $C^{(t)}(t\neq t^{*}$) in a neighborhood of $C_{1}$, local equigeneric deformation of each singular point of $C_{2}$, the fixed intersection multiplicity $2l_{1}+2l_{2}$ of $C^{(t)}$ and ${\operatorname{Tor}}(\sigma)$ at $w$, and intersection multiplicities of $C^{(t)}$ with ${\operatorname{Tor}}(\partial P_{\Delta})\setminus\{w\}$ prescribed by the condition $C^{(t)}\in{\mathcal{M}}_{0,n}(\Delta,{\boldsymbol{w}}(t))$. Since this deformation does not produce elliptic nodes in a neighborhood of $C_{1}$, the functions $W_{0}^{\kappa}(\widehat{\boldsymbol{w}}(t))$ and $\widetilde{W}_{0}^{\kappa}(\widehat{\boldsymbol{w}}(t)$ remain constant for $t\in(t^{*}-{\varepsilon},t^{*}+{\varepsilon})$.
[…]
Figure 5: Proof of Lemma 4.3, part (5), III
A sufficient condition for the existence of the required deformation can be written as
$h^{1}(C,{\mathcal{J}}_{Z/C}\otimes_{{\mathcal{O}}_{C}}{\operatorname{in}}^{*}{\mathcal{L}}_{P_{\Delta}})=0,$ (27)
where $C$ is the non-reduced union of a $2l_{1}$-multiple component $C_{1}$ and of $C_{2}$, and the zero-dimensional scheme $Z\subset C$ is supported at ${\operatorname{Sing}}\hskip 1.70709pt(C_{2})\cup(C\cap{\operatorname{Tor}}(\partial P_{\Delta}))\cup(C_{1}\cap C_{2})$ and is defined by the ideal $I^{cond}(C_{2}(z)$at each $z\in{\operatorname{Sing}}\hskip 1.70709pt(C_{2})$, by the ideal $I_{z}=\langle x^{2l_{1}},y-y_{i}\rangle$ at each point $(0,y_{i})\in C_{1}\cap C_{2}$, $i=1,...,N_{1}$, by the ideal
$\{\varphi\in{\mathcal{O}}_{C_{2},w_{i^{\prime}}^{\sigma^{\prime}}}\ :\ (\varphi\cdot C_{2})_{w_{i^{\prime}}^{\sigma^{\prime}}}\geq 2k_{i^{\prime}}^{\sigma^{\prime}}-1\}$
at some point $w_{i^{\prime}}^{\sigma^{\prime}}\in(C_{2}\cap{\operatorname{Tor}}(\partial P_{\Delta})))\setminus\{w\}$, by the ideal
$\{\varphi\in{\mathcal{O}}_{C_{2},w_{i^{\prime\prime}}^{\sigma^{\prime\prime}}}\ :\ (\varphi\cdot C_{2})_{w_{i^{\prime\prime}}^{\sigma^{\prime\prime}}}\geq 2k_{i^{\prime\prime}}^{\sigma^{\prime\prime}}\}$
at each point $w_{i^{\prime\prime}}^{\sigma^{\prime\prime}}\in C_{2}\cap{\operatorname{Tor}}(\partial P_{\Delta})\setminus\{w,w_{i^{\prime}}^{\sigma^{\prime}}\}$, and finally at the point $w$, by the ideal $I_{w}$ determined by the vanishing of the coefficients of the monomials $x^{r}$, $0\leq r<2l_{1}+2l_{2}$, $x^{r}y$, $0\leq r<2l_{1}$, see Figure 5. We establish (27) using the basic Horace method [13] (see also [11, Section 3.4.1]): the vanishing relation (27) follows from the series of vanishing relations
$\begin{cases}&h^{1}(C_{1},{\mathcal{J}}_{(Z:jC_{1})\cap C_{1}/C_{1}}\otimes_{{\mathcal{O}}_{C_{1}}}{\operatorname{in}}^{*}{\mathcal{L}}_{P_{\Delta}}(-jC_{1}))=0,\quad j=0,...,2l_{1}-1,\\ &h^{1}(C_{2},{\mathcal{J}}_{Z:(2l_{1}C_{1})/C_{2}}\otimes{{\mathcal{O}}_{C_{2}}}{\operatorname{in}}^{*}{\mathcal{L}}_{P_{\Delta}}(-2l_{1}C_{1}))=0.\end{cases}$ (28)
Each of the $2l_{1}$ relations in the upper line of (28) reads
$h^{1}(C_{1},{\mathcal{O}}_{C_{1}}(-w))=0,$
which follows by the Riemann-Roch theorem. The residue scheme $Z:(2l_{1}C_{1})$ in the lower line of (28) coincides with $Z$ at ${\operatorname{Sing}}\hskip 1.70709pt(C_{2})\cup(C_{2}\cap{\operatorname{Tor}}(\partial P_{\Delta}))\setminus\{w\}$ and is defined at $w$ by the ideal $I_{w}=\langle x^{2l_{2}},y\rangle$. Then, it lifts to $\widehat{C}_{2}\simeq{\mathbb{P}}^{1}$ in the form
$h^{1}(\widehat{C}_{2},{\mathcal{O}}_{\widehat{C}_{2}}({\boldsymbol{d}}))=0,$
where
$\deg{\boldsymbol{d}}=C_{2}^{2}-(C^{2}_{2}+C_{2}K_{{\operatorname{Tor}}(P_{\Delta})}+2)-(-C_{2}K_{{\operatorname{Tor}}(P_{\Delta})}-1)$
$=-1>-2,$
and hence (28) together with (27) follow.
(6) Suppose that $\xi$ is as in Lemma iv). Without loss of generality we can suppose that
@@
and that ${\Phi}$ is ramified at the points of $\widehat{\boldsymbol{w}}(t^{*})$ on the toric divisors associated with the segments $[(0,2m),(2p,0)]$ and $[(0,2m),(2q,0)]$, while the two points $w_{1}^{\sigma}(t),w_{2}^{\sigma}(t)\in{\boldsymbol{w}}(t)$ on the toric divisor ${\operatorname{Tor}}(\sigma)$, where $\sigma=[(2p,0),(2q,0)]$, merge to one point as $t\to t^{*}$. Then, $C={\Phi}({\mathbb{P}}^{1})$ admits a parametrization
@@
This deformation is regularly parameterized by the difference of the $x$-coordinates of $w_{2}^{\sigma}(t)$ and $w_{1}^{\sigma}(t)$ equal to $2ma\lambda+O(\lambda^{2})$. It follows from formulas (29) that the corresponding element $\xi(t)\in\overline{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))$ has the same orientations at a point on the toric divisor ${\operatorname{Tor}}([(0,2m),(2p,0)])$ for $\lambda<0$ and for $\lambda>0$, and so does for the point on the toric divisor ${\operatorname{Tor}}([(0,2m),(2q,0)])$, while the local branches at $w_{1}^{\sigma}(t)$, $w_{2}^{\sigma}(t)$ have opposite orientations with respect to the orientation of ${\operatorname{Tor}}_{\mathbb{R}}(\sigma)$. Furthermore, $\xi(t)$ has exactly two elliptic nodes in a neighborhood of each elliptic node of $C$. Thus, the constancy of $W_{0}^{\kappa}(t)$ and $\widetilde{W}_{0}^{\kappa}(t)$ follows. $\Box$
@@
Take two sequences $\widehat{\boldsymbol{w}}(0)=(\widehat{\boldsymbol{w}}_{\partial}(0),w_{0}(0))$ and $\widehat{\boldsymbol{w}}(1)=(\widehat{\boldsymbol{w}}_{\partial}(1),w_{0}(1))$, satisfying the conditions of Theorem 3.6. We may assume that $\widehat{\boldsymbol{w}}_{\partial}(0),\widehat{\boldsymbol{w}}_{\partial}(1)\in\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)$ for some $\rho>0$. Then, we join these sequences by a generic path $\{\widehat{\boldsymbol{w}}_{\partial}(t)\}_{0\leq t\leq 1}$, in $\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)$ and the points $w_{0}(0)$, $w_{0}(1)$ by a generic path $\{w_{0}(t)\}_{0\leq t\leq 1}$ in ${\mathcal{Q}}$, and verify the constancy of $W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ and $\widetilde{W}_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ in all possible wall-crossing events, for all quantum indices $\kappa$.